The correlation between parental control and student' self-esteem in a sample of 713 students was r = −0.19. the calculated t statistic was -4.178 for testing null hypothesis

Calculate an approximate P-value for testing H0 = 0 over the two-sided alternative .
Answer #1

Alpha = 0.05

df = n-2 = 713-2 = 711

alpha = 0.05

Null and Alternate Hypothesis

H0: ρ = 0 (No Correlation)

Ha: ρ <> 0 (Correlation is present)

Test Statistic

t = -4.178

p-value = TDIST(4.178,711,2) = 0.000033

Result

Since the p-value is less than 0.05, we reject the null hypothesis.